# Video Is Worth a Thousand Images: Exploring the Latest Trends in Long Video Generation

### v1

@@ 2.2.1. Autoencoder Formulation
Autoencoder is an unsupervised neural network that compresses its input to a compact latent layer and then learns to reproduce its input via backpropagation. The autoencoder is trained to minimize the reconstruction loss between the input $\mathbf{x}$ and the reconstructed output $\mathbf{\hat{x}}$. The most common application of autoencoder for video and long video generation is the construction of compressed latent space. For example, in  [84], authors use an encoder and decoder to project images from pixel to latent space to decrease the computational complexity of learning image distribution.
Variational Autoencoders (VAEs)  [47] are probabilistic generative models based on the principles of Bayesian networks. VAEs utilize a reconstruction strategy, where they project input video data from a high-dimensional space to a low-dimensional space and then reconstruct the video from this compact representation, also called the bottleneck layer, back to the original high-dimensional space. While traditional autoencoders learn compressed latent vectors, they cannot generate new data points, as they learn a fixed latent representation of the input data. VAEs solve this issue by converting the input into a distribution in the latent space rather than discrete points. An application of Variational Autoencoders (VAEs) for video generation is VQ-VAE [117]. VQ-VAE learns downstream latent representations using the Vector Quantized-Variational AutoEncoder (VQ-VAE), which is used in many video and image generation pipelines, such as VideoGen [152], VQGAN [20], and DALL-E [80]. In addition to using VAE models for compression, there is related work on using Variational Autoencoders for video generation. For instance, Hierarchical Patch VAE-GAN [24] uses a hierarchical approach to leverage both VAE and GAN techniques. VAEs have also been applied in video anomaly detection. In [131], the authors designed an LSTM-Convolutional-based Autoencoder to learn the distribution of single-scene videos, where the camera is fixed on a single background.
@@ 2.2.2. Masked Autoencoders
Masked autoencoder  [25] are scalable self-supervised learners for computer vision and have been used as the backbone for video generation by masking random patches of the input image and reconstructing the missing pixels. Video mask autoencoder is based on image mask autoencoder. During pre-training, a sizeable random portion of image patches are masked. The encoder processes only the remaining visible patches. After encoding, mask tokens are introduced, and the combined set of encoded patches and mask tokens is passed through a small decoder to reconstruct the original image at the pixel level. An extension of this approach to the video domain but using convolution instead of transformers is VideoMAC  [73]. It employs Masked Encoders and resource-friendly convNet architecture. VideoMAC is designed to reconstruct masked patches, randomly and symmetrically applied to frame pairs with a high masking ratio (0.75, as used in this study). Another notable model is the MAGVIT Masked Generative Video Transformer  [142]. MAGVIT outperforms existing diffusion and autoregressive models by order of magnitude in inference time. MAGVLT  [46] is also based on a masked autoencoder backbone. It proposed a unified generative vision-and-language (VL) model to produce images and text sequences.
@@ 2.3. Transformer Based Video Generation
GANs have limitations like mode collapse  [32], training instability, which needs fine-tuning of parameters, and a lot of training time and resources. Transformers, introduced in 2017  [118], made inroads into image and generation via autoregressive and masked encoding. Some of the key concepts to understand are Vision Transformers and Video Transformers.
@@ 2.3.1. Transformer Based Image Generation
Vision Transformers (ViT)  [18] represent a significant shift in the landscape of computer vision. The ViT architecture operates similarly to transformers used in natural language processing by dividing an image into smaller patches, like breaking text into tokens. DALL-E  [81] trained autoregressive transformer using 250 million image text pairs. They used ’Discrete VAE’  [83] to compress 256*256 images to 32 × 32 grid of image tokens. They concat these tokens with corresponding BPE text tokens and autoregressively learn the joint distribution of text and image tokens using a GPT transformer. CogView  [15] based on GPT transformer architecture outperforms DALL-E using FID distance, but DALL-E has better rendering capabilities of complex prompts. Autoregressive approaches like DALL-E and CogView suffer from slow generation because of their unidirectional token-by-token generation nature. CogView2  [17] improved the limitation of previous methods by training the Cross-Modal General Language Model by learning the joint distribution of image and text tokens using masking.
@@ 2.3.2. Autoregressive Based Video Generation
Video Transformers (VViT)  [3] represent a significant shift in the landscape of computer video generation. The VViT architecture functions like vision transformers  [18], but instead of tokenizing image patches, it tokenizes video patches. Phenaki [121] can generate arbitrary long videos conditioned on a sequence of prompts. Phenaki uses a pre-trained language model t5x for text embeddings,  [82]. It uses a variation of ViViT  [3], C-ViViT, which extracts and compresses video tokens to a minimum. It masked some tokens and learned to predict masked video tokens by the bidirectional transformer. During inference, it can predict the next token using an autoregressive transformer. The compression mechanism allows us to train and predict very long video sequences. CogVideo  [35] extend CogView2  [17]. It uses hierarchical training with multiple frame rates to enhance the alignment of text-clip pairs, significantly improving generation accuracy, especially for movements with complex semantics. In its two-stage architecture for video generation, it first generates keyframes auto-regressive, and then another transformer interpolates between these frames.
@@ 2.4.1. Text To Image Feature Representation
The core principle behind text-based visual generation tasks is effectively pairing text with the visual content. Many visual generation pipelines leverage pre-existing image-text pair models like CLIP (Contrastive Language-Image Pretraining)  [76]. CLIP has been pre-trained using a contrastive learning approach that optimizes the cosine similarity between image and text embedding. Given CLIP’s robust performance, many visual generation models like DALL·E 2  [79] incorporate CLIP’s text embedding to leverage its superior semantic understanding. It allows these models to enhance their ability to generate visually relevant and contextually appropriate content, effectively bridging the gap between text and visual representation.
@@ 2.4.2. LLMs Based Video Guidance
Many visual generation models like LLM Director  [156] leveraged standalone large language models (LLM)  [6] to enhance their performance. By integrating LLM, visual generation models can benefit from advanced natural language processing capabilities, allowing them to interpret and generate more nuanced and contextually relevant descriptions of captions in a single or multiple prompt with detailed scenes. One example of this design is LLM grounded VDM  [52]. When paired with visual inputs, LLMs can transform simple image descriptions into more elaborate storytelling, adding layers of meaning and context that enhance the viewer’s experience. LLM can also act as director of the whole video generation process and create coherent script shown by Vlogger [158]. The details on how the recent long video generation methods leverage LLMs are explained in 3.2.1.
@@ 2.5. Diffusion Models
Diffusion models are now state-of-the-art in image and video diffusion. Diffusion-based image and video pipelines use many building blocks, such as variational autoencoders, transformers, and language understanding, and have introduced new baselines. We will briefly cover the history and progress of diffusion-based architecture. The foundational work  [92] established a crucial framework for this approach, applying principles from nonequilibrium thermodynamics to unsupervised learning. This insight paved the way for subsequent advancements in generative modeling. A probabilistic diffusion model  [30] is a parameterized Markov chain trained via variational inference to generate samples that closely resemble the data over a finite duration. These ideas are applied to image and video generation, as discussed in the next section.
@@ 2.5.1. Image Diffusion
Image diffusion models generate images by denoising noisy inputs through noise prediction.  [94] introduced a novel perspective by estimating the data distribution’s gradients, further enriching the understanding of generative processes. However, the significant breakthrough in this area came in with  [30] influential paper on ”Denoising Diffusion Probabilistic Models.” do forward and reverse diffusion processes in latent space. The general mechanism for conditioning from single modality like text is to have a conditional denoising decoder using a cross attention layer of transformer in UNET backbone trained with embedding from text encoder like CLIP  [76]. In addition to CLIP, text encoders like BERT  [13] or T5 Encoder  [78] can be used. In their paper [84], the authors implement this paradigm by conditioning image generation through modifications to the attention layer of the U-Net. They achieve this by concatenating embeddings from different modalities, such as text or others [118].
@@ 2.5.2. Video Generation From Diffusion models
Video diffusion models are built on the same diffusion architecture as image diffusion models, but video diffusion models also need to ensure consistency between frames. Image diffusion model can generate videos by either extending it into the temporal domain, as done in the Video Diffusion Model [31], or by generating keyframes using an image diffusion model with a modified attention and noise sampling strategy to maintain frame consistency. We can interpolate these keyframes to achieve smoother transitions, as discussed in FreeBloom. [38].
U-Net and Transformers are the two primary backbone architectures driving video generation models. Video Diffusion Models adopt a standard U-Net-based diffusion model setup but modify the architecture for video modeling. It extends the traditional 2D U-Net into a 3D architecture, where each feature represents a 4D tensor that includes frames, height, width, and channels. This 3D U-Net is factorized over space and time, incorporating spatial and temporal attention mechanisms. The spatial attention block focuses on relevant regions within a single frame, while the temporal attention block captures dependencies across different frames over the temporal domain. On the other hand, Sora [109] is based on a Transformer-based diffusion model [72], which operates on spacetime patches of video and image latent codes. Visual input is represented as a sequence of spacetime patches, which serve as input tokens for the Transformer, as explained in  [109]. Long video generation strategies are discussed in detail in the next section.
@@ 3. Long Video: Generation Paradigm
We summarized various video generation approaches into three core paradigms.
@@
  • Divide and Conquer Approach: Videos are produced by creating keyframes or short video segments guided by storyline prompts, often with the aid of a large language model (LLM).
@@ 3.1. Auto Regressive Approaches
The Autoregressive generation paradigm works by future frame prediction. It trains and performs inference of the model on previous video frames. Each frame serves as the input for the future frames. This prediction process ensures the generated videos are consistent and coherently structured. It generates a context from anchor frame (generated via text prompt), text-to-image embedding, image, or multimodal input [ [14][120], [ [113]]. It uses context and previous N frames and generates future N frames optionally based on additional prompts and prior frames. It ensures robust consistency between frames due to the sequential nature of frame transition. This is illustarted in Fig. 5. Nevertheless, it needs to improve performance bottlenecks when generating long videos, as it has to create frames sequentially and can not leverage parallelization.
CogVideo  [35] is one of the first open-source autoregressive transformers for long video generation. CogVideo builds on the pre-trained text-to-image model CogView2  [16] to inherit the knowledge gained from text-to-image pretraining. CogVideo has limitations on the length of the input sequence and the spatial resolution of frames (default 160 x 160 can be upsampled to 480 x 480). NUWA-Infinity [54] improved spatial resolution by adopting a hierarchical video generation model. One limitation of CogVideo and NUWA-Infinity is that they consume a single prompt as input. Phenaki  [120] can handle multiple progressive text prompts. The model manages to preserve the temporal coherence of the video while adapting to the new prompt. It uses a single-level autoregressive model instead of the two-level architecture and compresses dynamic tokens using C-ViViT (build on top of ViViT  [3] ). The compression mechanism allows it to train and predict very long video sequences. Phenaki uses a pre-trained language model t5x [82] for text embeddings. It masked some tokens and learned to predict masked video tokens by the bidirectional transformer. During inference, It can predict the next token using an autoregressive transformer. One of the limitations of Phenaki is the memory requirements for training. MeBT  [141] proposed a memory-efficient bidirectional transformer that overcomes the memory limitations of auto-regressive methods by proposing an efficient transformer for video synthesis that can fully leverage the long-range dependency of video frames during training, while being able to achieve fast generation with linear time complexity. MeBT employs an encoder-decoder architecture based on a fixed number of latent bottleneck tokens, and that fixed scale of latent tokens gives its memory efficiency.
@@
All approaches discussed [ [54]],[ [14]],[ [120]] are based on transformers. Grid Diffusion  [49] is based on diffusion. Grid Diffusion first used compression and represented video using an image created from keyframes, which covers the primary motions or events of the video. It is called a ’grid image,’ which consists of 4 subframes representing video keyframes. During the training phase, they masked these frames and learned to produce masked frames conditioned on previous grid-image and nonmasked images. This design paradigm is illustrated in Fig. 6. As they replaced the challenge of video generation with image generation, they can create long videos up to 128 frames autoregressively with high image quality (low FVD scores  [115]. They utilized transfer learning of pre-trained stable diffusion model. [84] and have used only 2 Nvidia A-100 for training. Fig. 6 explains this architecture.
@@
Previously described approaches [ [54],  [14][120][49]] have limitation to render complex compositional prompts. Compositional prompts describe spatio-temporal entities that have dynamic interaction. For example, let’s take the example of a prompt: a man is walking with a black dog on the right side, and a blue car is driving from the opposite side. VideoTetris  [113] take on the challenge of compositional prompts. VideoTetris addresses this issue by manipulating and composing the attention maps of spatially and temporally denoising networks. It supports long video generation with progressive compositional prompts, where ’progressive’ refers to continuous changes in objects’ position, quantity, and attributes by introducing spatiotemporal Compositional Diffusion, which manipulates cross-attention values in denoising networks to synthesize videos that adhere to complex or evolving instructions.
@@
Autautoregressive long video generation approaches have evolved from single prompt, low-resolution video of CogVideo  [14] animating single scene to modern architecture like VideoTetris, which can model complex science dynamics with multiple prompts and can maintain high image quality. Autoregressive approaches provide smooth motion transitions between frames but suffer from slow generation due to their sequential nature. Also, it offers less control over complex scene dynamics like actors, bounding boxes, and spatial relationships between entities. Divide and conquer approaches, as discussed in 3.2, attempt to solve this problem by generating parallel frames and, often with the help of LLM, create video blueprints that are better suited to handle complex dynamic scenes, actors, and spatial relationships. Table LABEL:table:autoregressivecatalog lists key papers with critical insights.
@@ 3.2. Divide And Conquer Paradigms
The basic theme of divide and conquer is to generate keyframes or short clips based on single or multiple prompts and then interpolate between these frames or clips. The system often uses an anchor image as a reference frame for generating all subsequent frames. The system generates each key frame independently, allowing for parallel processing. Some challenges with the divide and conquer approach are keeping semantic consistency, smooth motion transformation between interpolated frames, and frame quality. One key theme of the Divide and Conquer approach is separating the planning and video generation stages, differentiating it from the autoregressive approach as illustrated in Fig. 8. Divide And Conquer Paradigms can be divided into three sub-paradigms: the Large language model as director, the Intermediate transition model, and the Agent-based framework. Some milestone papers with timelines are illustrated in Fig. 7
Figure 7. Divide-and-conquer-timeline: We used the date these papers were published in online resources like arXiv or https://openreview.net/. Papers catalog here are Align your Latents [5], Gen-L-Video [124], Free-Bloom [37], VideoDirectorGPT [55], LLM-grounded Video Diffusion Models [52], SEINE [11], FlowZero [62], Mora [145], Vlogger [158], Vidgen [86], DreamFactory [135] and Kubrik [26]

### v2

@@ 2.2.1. Autoencoder Formulation
Autoencoder is an unsupervised neural network that compresses its input into a compact latent layer and then learns to reproduce its input through backpropagation. The autoencoder is trained to minimize the reconstruction loss between the input $\mathbf{x}$ and the reconstructed output $\mathbf{\hat{x}}$. The most common application of an autoencoder for long video and video generation is the construction of compressed latent space. For example, in [45], the authors use an encoder and decoder to project images from pixel space to latent space, thereby decreasing the computational complexity of learning the image distribution.
Variational Autoencoders (VAE)  [43] address the limitations of traditional autoencoders by learning latent distributions instead of fixed representations, allowing new data generation. The VQ-VAE variant  [46] has become foundational for video/image generation pipelines like VideoGen  [47], VQGAN  [48], and DALL-E  [49]. Hybrid approaches like Hierarchical Patch VAE-GAN  [50] combine VAEs with GANs, while applications extend to anomaly detection through architectures like LSTM-Convolutional VAEs  [51].
@@ 2.2.2. Masked Autoencoders
Masked autoencoders  [44] serve as scalable self-supervised backbones for video generation by reconstructing randomly masked image patches. The approach extends to video through models like VideoMAC  [52], which uses convolutional networks to reconstruct symmetrically masked frame pairs at high masking ratios (0.75). The framework has evolved into advanced variants, such as MAGVIT  [53], which achieves significantly faster inference than diffusion models, and MAGVLT  [54], which unifies vision-language generation under this paradigm.
@@ 2.3. Transformer Based Video Generation
GANs have limitations, such as mode collapse [55], training instability, which requires fine-tuning of parameters, and a significant amount of training time and resources. Transformers, introduced in 2017  [56], made inroads into image and generation via autoregressive and masked encoding. Some of the key concepts to understand are Vision Transformers and Video Transformers.
@@ 2.3.1. Transformer Based Image Generation
Vision Transformers (ViT)  [57] revolutionized computer vision by processing images as patch-based tokens, similar to NLP transformers. DALL-E  [58] pioneered this approach for image generation, using a Discrete VAE  [59] to compress images into 32×32 tokens and training a GPT-style transformer on 250M image-text pairs. CogView  [60] later surpassed DALL-E in FID scores but maintained weaker complex prompt rendering. Both autoregressive models suffered from slow generation due to token-by-token processing, a limitation addressed by CogView2  [61] through masked cross-modal training.
@@ 2.3.2. Autoregressive Based Video Generation
Video transformers (VViT) [62] extend vision transformers [57] by tokenizing video patches. Phenaki [63] generates long videos from text prompts using T5X embeddings [64] and C-ViViT, a variant of VViT [62] that compresses tokens and employs masked and autoregressive prediction for long sequences. CogVideo [14] builds on CogView2 [61], using hierarchical training for better text-video alignment and a two-stage process involving keyframe generation and interpolation.
@@ 2.4.1. Text To Image Feature Representation
The core principle behind text-based visual generation tasks is effectively pairing text with the visual content. Many visual generation pipelines leverage pre-existing image-text pair models, such as CLIP (Contrastive Language-Image Pretraining)  [26]. CLIP has been pre-trained using a contrastive learning approach that optimizes the cosine similarity between image and text embedding. Given CLIP’s robust performance, many visual generation models, such as DALL·E 2  [7], incorporate CLIP’s text embedding to leverage its superior semantic understanding. It allows these models to enhance their ability to generate visually relevant and contextually appropriate content, effectively bridging the gap between text and visual representation.
@@ 2.4.2. LLMs Based Video Guidance
Many visual generation models, such as LLM Director  [65], leverage standalone large language models (LLMs)  [66] to enhance their performance. By integrating LLM, visual generation models can benefit from advanced natural language processing capabilities, enabling them to interpret and generate more nuanced and contextually relevant descriptions of captions in single or multiple prompts, along with detailed scenes. One example of this design is LLM grounded VDM  [67]. When paired with visual inputs, LLMs can transform simple image descriptions into more elaborate storytelling, adding layers of meaning and context that enhance the viewer’s experience. LLM can also act as the director of the entire video generation process and create a coherent script, as shown by Vlogger [68]. The details on how the recent long video generation methods leverage LLMs are explained in 3.2.1.
@@ 2.5. Diffusion Models
Diffusion models have emerged as the state-of-the-art approach for image and video generation, combining components such as variational autoencoders, transformers, and language models. The foundational work  [69] established key principles by applying non-equilibrium thermodynamics to unsupervised learning, while  [70] advanced the field through parameterized Markov chains trained by variational inference. These breakthroughs created the basis for modern diffusion architectures in visual generation tasks.
@@ 2.5.1. Image Diffusion
Image diffusion models generate images through iterative denoising, with  [71] contributing gradient-based estimation methods and  [70] establishing the foundational DDPM framework. For a text-conditioned generation, models typically employ a U-Net with cross-attention layers using embeddings from CLIP  [26], BERT  [72], or T5  [73]. Robin Rombach et al. implemented this in [45] through modified attention layers that combine multimodal embeddings  [56].
@@ 2.5.2. Video Generation From Diffusion models
Video generation models primarily use two architectures: 3D U-Nets and Transformers. The U-Net approach extends 2D diffusion models to handle 4D tensors (frames × height × width × channels) through factorized spatial-temporal attention, where spatial attention focuses on intra-frame regions and temporal attention captures inter-frame dependencies. Alternatively, Sora  [74] implements a Transformer-based diffusion model  [75] that processes videos as space-time patch tokens, as detailed in  [74]. These architectural approaches for long-video generation are further explored in subsequent sections.
@@ 3. Long Video: Generation Paradigm
We summarize various video generation approaches into three core paradigms.
@@
  • Divide-and-conquer approach: Videos are produced by creating keyframes or short video segments guided by storyline prompts, often with the aid of a large language model (LLM).
@@ 3.1. Auto Regressive Approaches
The autoregressive generation paradigm creates videos by sequentially predicting future frames from previous ones, ensuring temporal coherence through this frame-by-frame approach [ [76][15][77]]. While effective for maintaining consistency (Fig. 5), this method faces computational limitations for long videos due to its sequential nature. Key implementations include CogVideo  [14], which extends CogView2  [78] but is constrained by sequence length and resolution (160×160, upscalable to 480×480), and NUWA-Infinity  [79] which improves resolution through hierarchical generation. Phenaki  [15] advances the paradigm by handling multiple prompts through C-ViViT  [62] compression and T5X embeddings  [64], though its bidirectional training requires significant memory.
A paradigm shift emerged with VideoPoet [80], which demonstrated that multimodal training—combining text, images, and audio in a decoder-only transformer—could overcome these quality limitations. By pretraining on diverse objectives and fine-tuning for specific tasks, VideoPoet achieved state-of-the-art zero-shot generation, proving that autoregressive models can produce high-fidelity videos when augmented with rich multimodal signals.
@@
All approaches discussed [ [79]],[ [76]],[ [15]] are based on transformers. Grid Diffusion  [82] is based on diffusion. Grid Diffusion first used compression and represented video using an image created from keyframes, which covers the primary motions or events of the video. It is called a ’grid image,’ which consists of 4 subframes representing video keyframes. During the training phase, they masked these frames and learned to produce masked frames conditioned on previous grid images and non-masked images. This design paradigm is illustrated in Fig. 6. As they replaced the challenge of video generation with image generation, they can create long videos up to 128 frames autoregressively with high image quality (low FVD scores  [83]. They utilized transfer learning from a pre-trained stable diffusion model [45] and trained on only two Nvidia A100 GPUs. Fig. 6 explains this architecture. Building on this insight, ARLON  [84] combines the strengths of autoregressive transformers and diffusion models through its Asymmetric Diffusion Transformer (AsymmDiT) and latent VQ-VAE, achieving 128× compression (8× spatial + 6× temporal downsampling) in a 12-channel latent space. This hybrid approach maintains motion fidelity while enabling scalable, high-quality synthesis—effectively bridging diffusion models (Sec. 2.5) and token-based methods.
@@
Long video generation faces critical challenges in memory management and temporal coherence across extended sequences. ARLON  [84] addresses this via a training-free autoregressive inference method using pre-trained diffusion models. Its sliding-window queue mechanism processes frames with progressively increasing noise levels: fully denoised frames are removed from the head of the queue, while new noisy frames are added to the tail. This approach enables the unbounded generation of infinitely long videos without retraining, maintaining computational efficiency through a fixed-size queue that prevents memory overload.
Previous methods for video generation, including NUWA-Infinity  [79], CogView  [76], Phenaki  [15], and GRID  [82], face limitations in rendering complex compositional prompts that describe dynamic spatiotemporal interactions—such as ”a man walking with a black dog on his right while a blue car drives from the opposite direction.” VideoTetris [77] addresses this challenge by introducing spatiotemporal Compositional Diffusion, which manipulates cross-attention maps in denoising networks to synthesize videos that adhere to intricate or evolving instructions. This approach enables the generation of long videos with progressive compositional prompts, where ”progressive” refers to continuous changes in object positions, quantities, and attributes, ensuring precise alignment of interacting entities across space and time.
@@
VideoPoet  [80] An autoregressive LLM for high-quality synthesis from multi modal inputs June 2024
@@
Arlon [84] AR for long-range temporal guidance and DiT for high-fidelity synthesis Jan 2025
Autoregressive video generation has progressed from CogVideo  [76]’s single-prompt, low-resolution outputs to modern systems like VideoTetris, which can model complex scientific dynamics with multiple prompts while preserving quality. While autoregressive methods excel at smooth motion transitions, their sequential nature results in slow generation and limited control over complex scene elements (actors, bounding boxes, spatial relationships). Divide-and-conquer approaches (Sec. 3.2) address these limitations by enabling parallel frame generation and leveraging LLMs for structured video blueprints, improving the handling of dynamic scenes. Key papers and insights are cataloged in Table LABEL:table:autoregressivecatalog.
@@ 3.2. Divide And Conquer Paradigms
The basic theme of divide and conquer is to generate keyframes or short clips based on single or multiple prompts and then interpolate between these frames or clips. The system often uses an anchor image as a reference frame for generating all subsequent frames. The system generates each key frame independently, allowing for parallel processing. Some challenges associated with the divide-and-conquer approach include maintaining semantic consistency, ensuring smooth motion transformation between interpolated frames, and achieving high-quality frames. One key theme of the Divide and Conquer approach is separating the planning and video generation stages, differentiating it from the autoregressive approach as illustrated in Fig. 8. Divide and Conquer Paradigms can be divided into three sub-paradigms: the Large Language Model as Director, the Intermediate Transition Model, and the Agent-Based Framework. Some milestone papers with timelines are illustrated in Fig. 7.
Figure 7. Divide-and-conquer timeline: We used the dates these papers were published in online resources, such as arXiv or https://openreview.net/. Papers catalog here are Align your Latents [23], Gen-L-Video [16], Free-Bloom [21], VideoDirectorGPT [11], LLM-grounded Video Diffusion Models [67], SEINE [25], FlowZero [22], Mora [93], Vlogger [68], Vidgen [94], DreamFactory [95] and Kubrik [96]